OpenStudy (anonymous):

Max was solving the equation below and isn’t sure if his answer is correct. Explain to Max how he can check his answer and then help him identify any errors he made. Provide the correct solution in your explanation. 3x+6 8 = 7x−1 6 24 1 ( 3x+6 8 ) = 24 1 ( 7x−1 6 ) 3(3x+6) = 4(7x−1) 9x+18 = 28x−1 9x−9x+18 = 28x−9x−1 18 = 19x−1 18+1 = 19x−1+1 19 = 19x 1 = x help plz @creegan january123

OpenStudy (yolo_king):

3(3x+6) = 4(7x−1) 9x+18 = 28x−1<---- Here is the problem this negative one should be a -4

OpenStudy (anonymous):

thank you so much